# Trade-offs in Data Memorization via Strong Data Processing Inequalities

### v1

@@ C.3 Boolean memorization upper bounds
Here, we describe algorithms to complement the lower bound and establish Item 3 of Theorem 4.3. We split the proof into the two considered regimes $n\lesssim\sqrt{d}$ and $n\gtrsim\sqrt{d}\log d$.
Algorithm when $n\lesssim\sqrt{d}$
@@
Let $C_{sub}>0$ be a sufficiently large constant. On input $n$ i.i.d. samples $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, let $\hat{\theta}$ be the bit-wise majority vote of $X_{1:n}$. Our algorithm chooses $t=\min\left(d,\frac{Cd}{n}\right)$ and returns the hypothesis
$h(X):=\mathbbm{1}\left\{\langle{\hat{\theta}}^{[1,t]},{X^{[1,t]}}\rangle\geq\sqrt{2\log(200)t}\right\}.$
Here, recall that for a vector $v\in\mathbb{R}^{d}$, the notation $v^{[1,t]}=(v_{1},\dots,v_{t},0,\dots,0)$ denotes its projection onto the first $t$ coordinates embedded in $\mathbb{R}^{d}$.
We prove that $h$ is, with high probability, a good predictor. First, for $X\sim\mathcal{P}_{0}$, the same argument as in Section C.1 proves that $\Pr_{X\sim\mathcal{P}_{0}}[h(X)=1]\leq\frac{1}{200}$.
It remains to show that $h$ classifies most $X\sim\mathcal{P}_{\theta}$ correctly. Indeed, for every coordinate $j\in[d]$, we have
$\Pr[\hat{\theta}_{j}=\theta_{j}]\geq\frac{1}{2}+C^{\prime}\sqrt{n}\lambda$
for some constant $C^{\prime}$. Also,
$\Pr[X_{j}=\theta_{j}]=\frac{1}{2}+\lambda$
These two combined would imply that
$\mathop{\mathbb{E}}\langle\hat{\theta}^{[1,t]},X^{[1,t]}\rangle\geq C^{\prime}\lambda^{2}\sqrt{n}t=C^{\prime\prime}\sqrt{t}.$
Similarly as in Section C.1, we have
$\mathrm{Var}\left[\langle\hat{\theta}^{[1,t]},X^{[1,t]}\rangle\right]\leq t.$
Therefore, Chebyshev’s inequality gives
$\Pr[\langle\hat{\theta}^{[1,t]},X^{[1,t]}\rangle\leq((C^{\prime\prime})^{2}-\sqrt{200})\sqrt{t}]\leq\frac{1}{200}.$
By setting $C_{sub}$ to be large enough, we can guarantee that $(C^{\prime\prime})^{2}-\sqrt{200}\geq\sqrt{2\log(200)}$, which would consequently yield that $\Pr_{X\sim\mathcal{P}_{\theta}}[h(X)=0]\leq\frac{1}{200}$. This completes the accuracy analysis of $h$.
Regarding memorization, observe that $h$ can be described using $O(\frac{Cd}{n})$ bits. Hence, we have
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq I(h;X_{1:n})\leq H(h)\leq O\left(\frac{d}{n}\right),$
@@
We now provide an accurate algorithm for which $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P})=O(d^{2}\exp(-d/\sqrt{n}))$, and so if $n\geq\tilde{C}\sqrt{d}\log d$, then $\mathcal{A}$ learns an accurate model $h$ with nearly no excess memorization, particularly, such that $I(h;X_{1:n})=O(d^{-\tilde{C}/2})$.
Similarly to the previously described algorithm, given $n$ i.i.d. samples $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, let $\hat{\theta}$ be the bit-wise majority vote of $X_{1:n}$. Our algorithm returns the hypothesis
@@
Noting that this is the same predictor as before but with no projection, it is easy to verify that the accuracy analysis follows from the same calculation as in the previous case. We will therefore prove the memorization bound, which follows from different arguments.
By definition,
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\mathop{\mathbb{E}}_{\theta\sim\mathcal{U}(\{\pm 1\}^{d})}\left[I(h_{\hat{\theta}}\mid\theta;S\mid\theta)\right]\leq\mathop{\mathbb{E}}_{\theta\sim\mathcal{U}(\{\pm 1\}^{d})}[H(h_{\hat{\theta}}\mid\theta)]~{}.$
We will show that for any $\theta$, it holds that $H(h_{\hat{\theta}}\mid\theta)\lesssim d^{-\Omega(\tilde{C})}$, thus implying the claimed bound. Fix $\theta$, and let $\gamma=\Pr_{X_{1:n}}[h_{\hat{\theta}}\neq h_{\theta}]$ which is the probability that the bit-wise majority vote did not reconstruct $\theta$. By Hoeffding’s bound, for any coordinate $j\in[d]:\,\Pr_{X_{1:n}}[\hat{\theta}_{j}\neq\theta_{j}]\leq\exp(-n\lambda^{2}/2)$ so by union bounding we see that $\gamma\leq d\exp(-n\lambda^{2}/2)$.
Considering the Markov chain $\theta\to X_{1:n}\to h_{\hat{\theta}}$, Fano’s inequality shows that
$H(h_{\hat{\theta}}\mid\theta)$ $\leq H_{2}(\gamma)+\gamma\log(|\{\pm 1\}^{d}|-1)$
$<H_{2}(\gamma)+\gamma d$
$\leq 2\sqrt{\gamma}+\gamma d$
$\leq 2\sqrt{d}\exp(-n\lambda^{2}/4)+d^{2}\exp(-n\lambda^{2}/2)$
$\leq 3d^{2}\exp(-n\lambda^{2}/2)~{},$
where we used the easily verifiable numerical bound $H_{2}(\gamma)\leq 2\sqrt{\gamma}$. Overall we see that
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d^{2}\exp(-n\lambda^{2}/2))~{},$
which under our parameter regime $\lambda=C/d^{1/4}$, yields the claimed bound.
@@ D.1 Sparse Boolean sample complexity $(n=1)$
Recall that in $\mathbf{P}_{\mathrm{sB}}$, the learning problem is parameterized by $\theta=(S,y)$ where $S$ is a random subset including each $i\in[n]$ with probability $\nu\approx\frac{1}{d^{1/2}}$ and $y\sim\{\pm 1\}^{d}$ is uniformly at random. The distribution $\mathcal{P}_{\theta}$ is a product distribution $X$ on {±$\{\pm 1\}^{d}$ where $X_{S}=y_{S}$ with probability one and $X_{\overline{S}}$ is uniformly at random.
We claim that learning with a single samples is possible so long as the set $S$ satisfies that $|S|\geq\sqrt{4\log(200)d}$. First, note that this event happens with probability at least $1-\exp(-\Omega(\sqrt{d}))$ if we choose $\nu=\frac{C}{d^{1/2}}$ for a large enough constant $C$. Now, suppose we have $|S|\geq\sqrt{4\log(200)d}$ and let $X_{1}$ be the training sample. Consider the hypothesis
@@
We verify that $h$ has a low generalization error. To see this, we first observe that for $X\sim\mathcal{P}_{0}$, it holds $\mathop{\mathbb{E}}[\langle X_{1},X\rangle]=0$, and by Hoeffding’s bound, we further have $\Pr[\langle X_{1},X\rangle\geq\sqrt{2\log(200)d}]\leq 0.005$. On the other hand, conditioned on $\theta$, for every $X\sim\mathcal{P}_{\theta}$, we have that $X_{S}=(X_{1})_{S}$ and $X_{j}\sim\{\pm 1\}$ for every $j\not\in S$. In this case, it is easy to see that $\langle X,X_{1}\rangle$ is a random variable with mean $|S|\geq\sqrt{2\log(200)d}$ and standard deviation $O(\sqrt{d-|S|})$. Similarly to the proof for the case of $\mathbf{P}_{\mathrm{B}}$, we see that $\Pr_{X\sim\mathcal{P}_{\theta}}[\langle X,X_{1}\rangle\leq\sqrt{2\log(200)d}]\leq 0.005$. Overall, we see that the predictor $h$ has error at most $0.01$ as a classifier between $\mathcal{P}_{\theta}$ and $\mathcal{P}_{0}$, as desired.
@@ D.2 Sparse Boolean memorization lower bound
In this section, we establish Item 2 of Theorem 4.4. We will deviate slightly from the framework of Theorem 3.7. In particular, we will establish a lower bound on $I(A(X_{1:n});X_{1:n})$ via the test/train SDPI as in Theorem 3.7, however, we will not prove a data generation SDPI. Instead, we directly upper bound $I(X_{1:n};\theta)$ by the Shannon entropy of $\theta$, which is at most $O(\sqrt{d}\log(d))$. Then, a lower bound on $I(A(X_{1:n});X_{1:n}\mid\theta)$ follows easily.
Suppose $\theta\sim\Psi$ and $(X,X_{1:n})\sim\mathcal{P}_{\theta}^{n+1}$. We show that the pair $(X,X_{1:n})$ satisfy $\rho_{n}$-SDPI with $\rho_{n}=\Theta(\frac{2^{n}}{d^{1/2}})$. We begin with a basic observation: recall that the parameter $\theta$ consists of a pair $(S,x_{S})$ where each element is included in $S$ with probability $\nu=\Theta(d^{1/2})$. Here, conditioning on $X$ would not change the distribution of $S$: namely, it is still the case that each $i$ is in $S$ with probability $\nu$. Furthermore, if $i\in S$ then we know the $i$-th bits of the training samples (namely, $(X_{j,i})_{j=1}^{n}$) all agree with $X_{0,i}$. Otherwise, $(X_{j,i})_{j=1}^{n}$ is uniformly distributed in {±$\{\pm 1\}^{n}$.
With this observation in mind, we introduce a random variable $Z\sim\mathrm{Bin}_{\frac{1}{2}+\xi}(X)$ for some $\xi$ to be specified. Consider the following post-processing of $Z$ that produces sequence of samples $\tilde{X}_{1},\dots,\tilde{X}_{n}$:
@@
We argue that, for a proper choice of $\xi$, the joint distribution of $(X,\tilde{X}_{1:n})$ is identical to that of $(X,X_{1:n})$. Indeed, for each $i\in[d]$, the probability that all of $(X_{j,i})_{j=1}^{n}$ are the same is $\nu+(1-\nu)2^{-n+1}$, which is also the case for $\tilde{X}_{1:n}$. Furthermore, conditioning on the event that $(X_{j,i})_{j=1}^{n}$ are not all-00 nor all-$1$, they are uniformly distributed in {±$\{\pm 1\}^{n}\setminus\{(-1)^{n},(+1)^{n}\}$, which is, again, also the case for $\tilde{X}_{1:n}$.
We turn to the case that all of $(X_{j,i})_{j=1}^{n}$ are the same. It is easy to calculate that the probability of the event $(X_{j,i})_{j=1}^{n}=(X_{0,i})^{n}$ is $\nu+(1-\nu)2^{-n}$, while the same probability term w.r.t. $\tilde{X}$ evaluates to $(\frac{1}{2}+\xi)\cdot(\nu+(1-\nu)2^{-n+1})$. It remains to set $\xi$ properly so that the two probability quantities coincide. Indeed, this gives
@@
Recall that we set $\nu=\Theta(d^{-1/2})$. Therefore, for all $n\leq o(\log d)$, the expression simplifies to $\xi=\Theta(\nu 2^{n})$.
We have shown that with the proper choice of $\xi$, we can define $Z\sim\mathrm{Bin}_{\frac{1}{2}+\xi}(X)$ and post-process $Z$ to obtain a sequence of samples $\tilde{X}_{1:n}$ identically distributed as $X_{1:n}$. This implies that the pair $(X,X_{1:n})$ enjoys at least the same SDPI as the pair $(X,Z)$. Quantitatively, the SDPI parameter we achieve is $\rho_{n}=\xi^{2}=\Theta(\nu^{2}2^{2n})=\Theta(\frac{2^{2n}}{d})$. This consequently implies that
$I(\mathcal{A}(X_{1:n});X_{1:n})\geq\frac{1}{\rho_{n}}I(\mathcal{A}(X_{1:n});X)=\Omega\left(\frac{d}{2^{2n}}(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)\right),$
as claimed. For all $n\leq o(\log d)$, this lower bound is far above $\sqrt{d}\log d$. Hence,
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})=I(\mathcal{A}(X_{1:n});X_{1:n}\mid\theta)\geq I(\mathcal{A}(X_{1:n});X_{1:n})-H(\theta)=\Omega\left(\frac{d}{2^{2n}}\right).$

### v2

@@ C.3 Boolean memorization upper bounds
Here, we describe algorithms to complement the lower bound and establish Item 3 of Theorem 4.3. We split the proof into the two considered regimes $n\leq\sqrt{d}$ and $n\gtrsim\sqrt{d}\log d$.
Algorithm when $n\leq\sqrt{d}$
@@
Let $C_{sub}>0$ be a sufficiently large constant. On input $n$ i.i.d. samples $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, let $\hat{\theta}$ be the bit-wise majority vote of $X_{1:n}$. Our algorithm chooses $t=\min\left\{d,\frac{C_{sub}d}{n}\right\}$ and returns the hypothesis
$h(X):=\mathbbm{1}\left\{\langle{\hat{\theta}}^{[1:t]},{X^{[1:t]}}\rangle\geq\sqrt{2\log(200)t}\right\}.$
Here, recall that for a vector $v\in\mathbb{R}^{d}$, the notation $v^{[1:t]}=(v_{1},\dots,v_{t},0,\dots,0)$ denotes its projection onto the first $t$ coordinates embedded in $\mathbb{R}^{d}$.
We prove that $h$ is, with high probability, a good predictor. First, for $X\sim\mathcal{P}_{0}$ it holds that $\Pr_{X\sim\mathcal{P}_{0}}[h(X)=1]\leq\frac{1}{200}$ by Hoeffding’s inequality (as in Section C.1).
It remains to show that $h$ classifies most $X\sim\mathcal{P}_{\theta}$ correctly. This will follow from Slud’s inequality [Slud, 1977] paired with a standard Gaussian tail bound, as stated below.
Lemma C.5.
For any coordinate $j\in[d]$, the majority vote satisfies $\Pr[\hat{\theta}_{j}=\theta_{j}]\geq\frac{1}{2}+\sqrt{1-\exp(-2n\lambda^{2})}$.
In particular, a Taylor approximation ensures that $\Pr[\hat{\theta}_{j}=\theta_{j}]\geq\frac{1}{2}+C^{\prime}\sqrt{n}\lambda$ for some constant $C^{\prime}$. Also, $\Pr[X_{j}=\theta_{j}]=\frac{1}{2}+\lambda$. These two combined would imply that
$\mathop{\mathbb{E}}\left[\langle\hat{\theta}^{[1:t]},X^{[1:t]}\rangle\right]\geq C^{\prime}\lambda^{2}\sqrt{n}t\geq C^{2}\sqrt{t}~,$
the latter holding for sufficiently large absolute constant $C_{sub}$, under the assumption that $n\leq\sqrt{d}$. Note that the bound above is the same as Eq. (6), and the proof follows similarly by applying Hoeffding’s inequality as in Eq. (7)
To upper bound the excess memorization of $\mathcal{A}$, observe that $h$ can be described using $O(\frac{C_{sub}d}{n})$ bits. Hence, we have
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq I(h;X_{1:n})\leq H(h)=O\left(\frac{d}{n}\right),$
@@
We now provide an accurate algorithm for which $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d^{2}n\exp(-nC^{2}/(2\sqrt{d}))$. In particular, for which there exists a fixed constant $\tilde{C}$ such that for $n\geq\tilde{C}C^{2}\sqrt{d}\log d$, $\mathcal{A}$ learns an accurate model $h$ with nearly 0 excess memorization. Similarly to the previously described algorithm, given $n$ i.i.d. samples $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, let $\hat{\theta}$ be the bit-wise majority vote of $X_{1:n}$. Our algorithm returns the hypothesis
@@
Noting that this is the same predictor as before but with no projection, it is easy to verify that the accuracy analysis follows from a similar calculation as in the previous case. We now establish the memorization bound. By definition,
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=I(h_{\hat{\theta}};S\mid\theta)\leq H(h_{\hat{\theta}}\mid\theta)=H(\hat{\theta}\mid\theta)~.$
We will show that for any fixed setting $\theta^{\prime}\in\{\pm 1\}^{d}$, it holds that
$H(h_{\hat{\theta}}\mid\theta=\theta^{\prime})=O(d^{2}n\exp(-nC^{2}/(2\sqrt{d}))~,$
thus implying the claimed bound. Assume $\theta=\theta^{\prime}$, and let $\gamma=\Pr_{X_{1:n}}[\hat{\theta}\neq\theta^{\prime}]$ which is the probability that the bit-wise majority vote did not reconstruct $\theta$. By Hoeffding’s bound, for any coordinate $j\in[d]:\,\Pr_{X_{1:n}}[\hat{\theta}_{j}\neq\theta^{\prime}_{j}]\leq\exp(-n\lambda^{2}/2)$ so by union bounding we see that $\gamma\leq d\exp(-n\lambda^{2}/2)$.
This means that the distribution of $\hat{\theta}$ conditioned on $\theta=\theta^{\prime}$ has TV distance of at most $\gamma$ to the distribution where $\theta$ is fixed to $\theta^{\prime}$ (which has 0 entropy). Applying Lemma A.1,
$H(\hat{\theta}\mid\theta)$ $=\mathop{\mathbb{E}}_{\theta^{\prime}\sim\mathcal{U}(\{\pm 1\}^{d})}[H(\hat{\theta}\mid\theta=\theta^{\prime})]$
$\leq H(\theta\mid\theta=\theta^{\prime})+2\gamma\log(2^{d}/\gamma)$
$=2\log 2\cdot\gamma d+2\gamma\log(1/\gamma)$
$=\left(2\log 2\cdot d^{2}+dn\lambda^{2}\right)\exp(-n\lambda^{2}/2)$
$=O(d^{2}n\cdot\exp(-nC^{2}/(2\sqrt{d}))~,$
where we used $\lambda=C/d^{1/4}$.
@@ D.1 Sparse Boolean sample complexity $(n=1)$
The following sample complexity analysis is similar to that of Brown et al. [2021].
Recall that in $\mathbf{P}_{\mathrm{sB}}$, the learning problem is parameterized by $\theta=(S,y)$ where $S$ is a random subset including each $i\in[n]$ with probability $\nu=\frac{C}{\sqrt{d}}$ and $y\sim\mathcal{U}(\{\pm 1\}^{d})$ is chosen uniformly at random. The distribution $\mathcal{P}_{\theta}$ is a product distribution $X$ on $\{\pm 1\}^{d}$ where $X_{S}=y_{S}$ with probability one and $X_{\overline{S}}$ is chosen uniformly at random.
We claim that learning with a single sample is possible as long as the set $S$ satisfies that $|S|\geq\sqrt{4\log(200)d}$. First, note that this event happens with probability at least $1-\exp(-\Omega(\sqrt{d}))$ if we choose $\nu=\frac{C}{d^{1/2}}$ for a large enough constant $C$. Now, suppose we have $|S|\geq\sqrt{4\log(200)d}$ and let $X_{1}$ be the training sample. Consider the hypothesis
@@
We verify that $h$ has a low generalization error. To see this, we first observe that for $X\sim\mathcal{P}_{0}$, it holds that $\mathop{\mathbb{E}}[\langle X_{1},X\rangle]=0$, and by Hoeffding’s bound, we further have $\Pr[\langle X_{1},X\rangle\geq\sqrt{2\log(200)d}]\leq 0.005$. On the other hand, conditioned on $\theta$, for every $X\sim\mathcal{P}_{\theta}$, we have that $X_{S}=(X_{1})_{S}$ and $X_{j}\sim\mathcal{U}(\{\pm 1\})$ for every $j\not\in S$. In this case, it is easy to see that $\langle X,X_{1}\rangle$ is a random variable with mean $|S|\geq\sqrt{4\log(200)d}$ given by a sum of independent $\pm 1$’s. By Hoeffding bound, we get that $\Pr_{X\sim\mathcal{P}_{\theta}}[\langle X,X_{1}\rangle\leq\sqrt{2\log(200)d}]\leq 0.005$. Overall, we see that the predictor $h$ has the expected error of at most $0.01$.
@@ D.2 Sparse Boolean memorization lower bound
We now establish item 2 of Theorem 4.4. Our proof will deviate from the framework of Theorems 3.7 and 3.8. Instead, we will establish a test/train SDPI via a more direct argument that still however relies on our notion of dominating random variables. In addition, we will use a direct upper bound on the description length of $\theta$ to upper bound $I(X_{1:n};\theta)$ instead of relying on the data generation SDPI.
We will rely on the following simple version of our reduction to dominating variables in Theorem 3.8.
Lemma D.1.
Let $\mathbf{P}=(\mathcal{P}_{\theta})_{\theta\sim\Psi}$ be a learning problem, and suppose that there exists a random variable $Z_{X}^{\mathrm{train}}$ (that depends on the value of $X$) such that there is a mapping $\Phi^{\mathrm{train}}$ satisfying: the distribution of $(X,\Phi^{\mathrm{train}}(Z_{X}^{\mathrm{train}}))$ when $\theta\sim\Psi$ and $X\sim P_{\theta}$ is identical to the distribution of $(X,X_{1:n})$ when $\theta\sim\Psi$ and $X_{1:n}\sim P_{\theta}^{n}$. Then if the pair of random variables $(X,Z_{X}^{\mathrm{train}})$ satisfies $\rho_{n}$-SDPI, then $(X,X_{1:n})$ for $X\sim\mathcal{P}_{\theta}$, $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$ and $\theta\sim\Psi$ satisfies $\rho_{n}$-SDPI.
The proof of this lemma follows from the fact that if a pair of random variables $(A,B)$ satisfies $\rho$-SDPI, then for any (randomized) postprocessing function $\Phi$, the pair of random variables $(A,\Phi(B))$ also satisfies $\rho$-SDPI.
For completeness we also state a variant of Theorem 3.7 that bypasses the data generation SDPI (the proof is implicit in the proof of Thm. 3.7).
Theorem D.2.
Let $\mathbf{P}=(\mathcal{P}_{\theta})_{\theta\sim\Psi}$ be a learning problem satisfying the following test/train SDPI: the variables $(X,X_{1:n})$ for $\theta\sim\Psi$, $X\sim\mathcal{P}_{\theta}$ and $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$ satisfy $(\rho_{n},\delta_{n})$-SDPI. Then any algorithm $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{M}$ for $\mathbf{P}$ satisfies the excess memorization bound:
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P})\geq\frac{I(\mathcal{A}(X_{1:n});X)-8\delta_{n}\log(|\mathcal{M}|/\delta_{n})}{\rho_{n}}-I(\mathcal{A}(X_{1:n});\theta)\ .$
Moreover, for any $\alpha<\frac{1}{2}:$
$\mathrm{mem}_{n}(\mathbf{P},\alpha)\geq\frac{C_{\alpha}-8\delta_{n}\log(|\mathcal{M}|/\delta_{n})}{\rho_{n}}-I(\mathcal{A}(X_{1:n});\theta)~,\text{~~~where~~~}C_{\alpha}:=(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~.$
The test/train SDPI via post-processing.
Suppose $\theta\sim\Psi$ and $(X,X_{1:n})\sim\mathcal{P}_{\theta}^{n+1}$. We show that the pair $(X,X_{1:n})$ satisfies $\rho_{n}$-SDPI with $\rho_{n}=\Theta(2^{2n}/d)$.
We begin with a basic observation: recall that the parameter $\theta$ consists of a pair $(S,y_{S})$ where each element is included in $S$ with probability $\nu=\Theta\left(\sqrt{d}\right)$. Here, conditioning on $X$ would not change the distribution of $S$: namely, it is still the case that each $i$ is in $S$ with probability $\nu$. Furthermore, if $i\in S$ then we know the $i$-th bits of the training examples (namely, $X_{1,i},X_{2,i},\dots,X_{n,i}$) all agree with $X_{i}$ Otherwise, $(X_{1,i},\dots,X_{n,i})$ is uniformly distributed in $\{\pm 1\}^{n}$.
With this observation in mind, we introduce a random variable $Z_{X}^{\mathrm{train}}=Z\sim\mathrm{BSC}_{\frac{1}{2}-\xi}(X)$ for some $\xi$ to be specified. Consider the following post-processing $\Phi^{\mathrm{train}}(Z)$ that produces a sequence of example $\tilde{X}_{1},\dots,\tilde{X}_{n}$ as follows:
@@
The correctness of post-processing.
We compare the marginal distribution of $(X,\tilde{X}_{1:n})$ and $(X,X_{1:n})$. Let us prove that they are identically distributed for a proper choice of $\xi$. First of all, we should note that due to definition of $\theta,X,Z,X_{1:n}$ and $\Phi^{\mathrm{train}}$, we have that $(X,\tilde{X}_{1:n})$ and $(X,X_{1:n})$ are both product distribution over $(\{\pm 1\}^{1+n})^{d}$ (i.e., the $d$ coordinates of the test example and the $n$ examples can be understood as independently generated). Next, observe that for each $j\in[d]$, the probability that all of $(X_{1,j},\dots,X_{n,j})$ are the same is $\nu+(1-\nu)2^{-n+1}$, which is also the case for $\tilde{X}_{1:n}$. Furthermore, conditioning on the event that $(X_{j,i})_{j=1}^{n}$ are not all-0 nor all-$1$, they are uniformly distributed in $\{\pm 1\}^{n}\setminus\{(-1)^{n},(+1)^{n}\}$, which is, again, also the case for $\tilde{X}_{1:n}$.
We turn to the case that all of $X_{*,j}$ are the same. It is easy to calculate that the probability of the event $X_{*,j}=(X_{0,i})^{n}$ is $\nu+(1-\nu)2^{-n}$, while the same probability term w.r.t. $\tilde{X}$ evaluates to $(\frac{1}{2}+\xi)\cdot(\nu+(1-\nu)2^{-n+1})$. It remains to set $\xi$ properly so that the two probability quantities coincide. Indeed, this gives
@@
Recall that we set $\nu=C/\sqrt{d}$. Therefore, for $n\leq\log_{2}(\sqrt{d}/C)$, it holds that $\xi\leq\nu 2^{n}$.
We have shown that with the proper choice of $\xi$, we can define $Z\sim\mathrm{BSC}_{\frac{1}{2}-\xi}(X)$ and post-process $Z$ to obtain a sequence of samples $\tilde{X}_{1:n}$ identically distributed as $X_{1:n}$. By Lemma D.1, this implies that the pair $(X,X_{1:n})$ enjoys at least the same SDPI as the pair $(X,Z)$. Quantitatively, the SDPI parameter we achieve is $\rho_{n}=\xi^{2}\leq\nu^{2}2^{2n}=\frac{C^{2}2^{2n}}{d}$. By Theorem D.2, for a constant $\alpha<1/2$ this consequently implies that
$\mathrm{mem}_{n}(\mathbf{P},\alpha)\geq\frac{C_{\alpha}}{d}{C^{2}2^{2n}}-I(\mathcal{A}(X_{1:n});\theta)~.$
In particular, there exists a constant $c_{\alpha}$ such that for any $n\leq\log_{2}d/2-c_{\alpha}$ we have that $\frac{C_{\alpha}}{d}{C^{2}2^{2n}}\geq 2C\sqrt{d}\log d$. Noting that
$I(\mathcal{A}(X_{1:n});\theta)\leq H(\theta)=d(-(1-\nu)\log(1-\nu)-\nu(\log(\nu/2)))=(1+o(1))C\sqrt{d}\log(\sqrt{d}/C),$
we obtain
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})\geq\frac{C_{\alpha}}{d}{C^{2}2^{2n}}-H(\theta)=\Omega\left(\frac{d}{2^{2n}}\right).$